OpenStudy (anonymous):

3. The x-intercept of a line is –5 and the y-intercept of the line is –2. What is the equation of the line?

OpenStudy (anonymous):

A. y = -5/2x -5 B. y = 2/5x + 2 C. y = 5/2x - 2 D. y = -2/5x - 2

OpenStudy (anonymous):

Helpppp???